OpenStudy (anonymous):

What's the best first step to eliminate the y-value in the system of equations? x + 8y = –4 2x – 4y = –13

OpenStudy (anonymous):

System of Equations will be the death of me...

OpenStudy (anonymous):

How the hell does this make sense? Has anyone ever been far even as decided to use even go want to do look more like?

OpenStudy (dan815):

hey

OpenStudy (anonymous):

hai. :)

OpenStudy (dan815):

you can add equations together its no different from adding numbers together

OpenStudy (dan815):

some equation = -4 some equation = -13 so if i add -4-13 that is the same as adding those 2 equations together

OpenStudy (anonymous):

I know i need to multiply one of the two equations by 2 or -2.

OpenStudy (dan815):

that is the theory behind it

OpenStudy (dan815):

now u just want to add in a way so that one of those variables is eliminated

OpenStudy (jdoe0001):

notice the functions themselves $$ \begin{matrix} x& \color{blue}{+ 8y} &= &–4\\ 2x & \color{blue}{– 4y} &= &–13 \end{matrix} $$ to eliminate "y" you'd want to make the "y's" the same value but different sign so above you have +8y, so +8y -8y ----- 0 thus eliminated

OpenStudy (dan815):

-6*

OpenStudy (dan815):

x + 8y = –4 (2x – 4y = –13) *2 -->(4x – 8y = –26) add them x + 8y = –4 4x – 8y = –26 -------------- 5x = -30 x =-6

OpenStudy (anonymous):

i only need the first step, heres the review question im stuck on: Select the best possible first step to solving the system by first eliminating the y variable. x + 8y = –4 2x – 4y = –13 a. Multiply the first equation by 2. b. Multiply the second equation by 2. c. Multiply the first equation by –2. d. Multiply the second equation by –2.

OpenStudy (compassionate):

Hello, my name is Compassionate. But you already know that, ya? System of equations is fairly easy. There are two methods that are very easy to use. The first method is called the Elimination method and the second method is called the substitution method. They both have different usages. Foe example, in the substitution method, you SOLVE for either equation then substitute it into the other equation. For example: x = y + 3 x + y = 2 I can see here that x is already solved. So if I know what x equals then I just plug it into the equation. (y + 3) + y = 2 y + 2 + y = 2, now, I just solve for y. And once I solve for y -- I plug y into either equation, and there we go! I got my x and y points. Now the elimination method is different. We want to ELIMINATE a term, either x or y, then solve. In the equations above: x + 8y = –4 2x – 4y = –13 We can eliminate either one. But lets eliminate x first. So, what do you need to multiply EITHER of the equations by to get them to cancel x terms? Well, wee need to multiply the first one by NEGATIVE 2. -2(x + 8y = -4) 2x - 4y = - 13 -2x - 16y = 8 2x - 4y = - 13 Now, we see, upon adding, the xs will cancel. -20y = -5 Now we just solve like a normal equation (Because we eliminated x, we're just left to solve for y!) y = -20/-5 y = 4 (Remember, dividing a negative by a negative = positive) SO, now that you know that y = 4, take 4 and plug it into EITHER of the equations and solve it for X! :) You can do the rest. I believe in you!

OpenStudy (anonymous):

none of your answers have matched, or come close to, A.,B.,C., or D.

OpenStudy (anonymous):

C.! C is for Compassionate! Thank you! You went out of your way and solved the full damn thing too! Thanks!

OpenStudy (compassionate):

You're welcome ~~

OpenStudy (anonymous):

All I needed: We can eliminate either one. But lets eliminate x first. So, what do you need to multiply EITHER of the equations by to get them to cancel x terms? Well, wee need to multiply the first one by NEGATIVE 2.

OpenStudy (dan815):

lol xD no

OpenStudy (anonymous):

Thank you everybody for trying.

OpenStudy (compassionate):

Hold up there, Davis, I may have made a slight miscalculation

OpenStudy (anonymous):

could you try this one @Compassionate? Eliminate the x-value. 4x + 9y = 21 3x – 8y = 1 a. ultiply the first equation by 3 and multiply the second equation by 4. b. Multiply the first equation by –3 and multiply the second equation by –4. c. Multiply the first equation by 8 and multiply the second equation by 9. d. None of the above

OpenStudy (anonymous):

pellet.

OpenStudy (compassionate):

No problem, Davis, can you give me a minute to calculate something? Thanks ~

OpenStudy (compassionate):

Hey, sorry for the short wait. Would you mind browsing Facebook, messing around on YouTube? Or is this a timed test. I'm at BK and making an order currently. Can you wait a few minutes?

OpenStudy (anonymous):

it's like you know me perfectly, im on Facebook, and YouTube, and its just an entry-review.

OpenStudy (compassionate):

Alright. They sort of ran out of ice cream and whatnot. So I'm having to wait on that.

OpenStudy (compassionate):

(I'm talking about your original question) The first equation can be multiplied by 2 or negative two... it just seems easier to multiply it by 2 and solve for y. So, it should be, "Multiply the first one by 2" This one makes more sense. I admit, I had to look at it for awhile. While it doesn't matter if we multiply it by 2 or -2, doing so by 2 eliminates a single step.

OpenStudy (dan815):

it should be "multiply the 2nd one by 2"

OpenStudy (anonymous):

Okay, thank you for clearing that up with me, it's just these types of questions I dont understand.

OpenStudy (dan815):

the SECOND* okay

OpenStudy (dan815):

your original question wants you to eliminate y

OpenStudy (anonymous):

So B? and if you understand so well Dan, do the second question.

OpenStudy (anonymous):

While Compassionate enjoys some BK Ice Cream.

OpenStudy (compassionate):

Oops! Dan is right! I didn't read it correctly. It wants you to eliminate Y first. So multiply the second one by 2. SORRY! Read it wrong. But I hope you now understand how to solve system of equations!

OpenStudy (anonymous):

ehh, it needs more than one example for me to understand things. usually two of the same thing.

OpenStudy (dan815):

sure ill give you another example :)

OpenStudy (dan815):

3x + 2y =5 x-6y=3 eliminate y

OpenStudy (anonymous):

Eliminate the x-value. 4x + 9y = 21 3x – 8y = 1 a. ultiply the first equation by 3 and multiply the second equation by 4. b. Multiply the first equation by –3 and multiply the second equation by –4. c. Multiply the first equation by 8 and multiply the second equation by 9. d. None of the above

OpenStudy (anonymous):

mine has multiple choice, which is what i have to do for my review, and my test.

OpenStudy (dan815):

well do this one, 3x + 2y =5 x-6y=3 eliminate y

OpenStudy (dan815):

what will be your first step?

OpenStudy (dan815):

hello??

OpenStudy (dan815):

help me help you

OpenStudy (compassionate):

Listen to Dan, Davis, help us help you.

OpenStudy (anonymous):

y=mx+b form?

OpenStudy (anonymous):

i'm clueless...

OpenStudy (compassionate):

We want to eliminate x: 4x + 9y = 21 3x – 8y = 1 that is: 4x and 3x What do we need to multiply 4x and 3x by to get them BOTH to cancel?

OpenStudy (anonymous):

like i said, clueless..

OpenStudy (compassionate):

What is the lowest common multiple of 4 and 3?